# Estimating Causal Effects for Binary Outcomes Using Per-Decision Inverse Probability Weighting

### v2

@@ 5 Application
We apply the proposed pd-EMEE to the Drink Less MRT data. Recall that there are $n=349$ participants in this MRT, each in the study for $T=30$ days. Each participant was randomized once every day at 8pm with 0.6 probability to receive a push notification encouraging using the Drink Less app ($A_{t}=1$) and with 0.4 probability to receive nothing ($A_{t}=0$). Individuals were eligible to be randomized at all decision points ($I_{t}\equiv 1$). The proximal outcome, $Y_{t,\Delta}$, is whether the individual logged in to the app in the subsequent 72 hours, so $\Delta=3$. The binary sub-outcome, $R_{t+1}$, is whether the individual logged in to the app between 8 pm of day $t$ and 7:59 pm of day $t+1$, and $Y_{t,\Delta}=\max(R_{t+1},R_{t+2},R_{t+3})$.
We used pd-EMEE and EMEE for two estimands: the fully marginal effect and the effect moderated by $\text{Day}_{t}$, the number of days in the study. $\text{Day}_{t}$ takes values $0,1,\ldots,29$. The fully marginal excursion effect, $\beta_{0}$, is defined by setting $S_{t}=1$ in (5):
$\beta_{0}=\log\frac{E\left\{Y_{t,3}(\bar{A}_{t-1},1,\bar{0}_{2})\right\}}{E\left\{Y_{t,3}(\bar{A}_{t-1},0,\bar{0}_{2})\right\}}.$
The effect moderated by $\text{Day}_{t}$, parameterized by $(\beta_{1},\beta_{2})$, is defined by setting $S_{t}=(1,\text{Day}_{t})$ in (5):
$\beta_{1}+\beta_{2}\text{Day}_{t}=\log\frac{E\left\{Y_{t,3}(\bar{A}_{t-1},1,\bar{0}_{2})\mid\text{Day}_{t}\right\}}{E\left\{Y_{t,3}(\bar{A}_{t-1},0,\bar{0}_{2})\mid\text{Day}_{t}\right\}}.$
$\beta_{1}$ is the causal excursion effect on the first day, and $\beta_{2}$ is the change in the effect with each additional day. The $g(H_{t})$ in the working model for $E\{Y_{t,3}(\bar{A}_{t-1},0,\bar{0}_{2})\mid H_{t},I_{t}=1\}$ include gender, age, employment type, the AUDIT score (a measure of alcohol misuse), and two time-varying variables about app usage: whether the individual used the app before 8 pm that day, and whether they used the app after 9 pm the previous day.
For each of $\beta_{0},\beta_{1}$, and $\beta_{2}$, pd-EMEE and EMEE yield estimates in the same direction, and pd-EMEE results in smaller standard errors and narrower confidence intervals (Table 4). For example, $\beta_{0}$ estimated using pd-EMEE is 0.127 with a 95% confidence interval (0.075, 0.180). This means that when averaged over the course of the 30-day study, a notification (followed by 2 days with no notifications) increases the chance of someone logging into the app in the next 72 hours by approximately 13.5% ($=e^{0.127}-1$). The same $\beta_{0}$ estimated using EMEE is 0.079 with a 95% confidence interval (-0.031, 0.188), about twice as wide as that obtained by pd-EMEE. The 95% confidence intervals for $\beta_{1}$ and $\beta_{2}$ obtained using pd-EMEE are also substantially narrower than those obtained using EMEE. This illustrates the substantial efficiency improvement by the proposed pd-EMEE over the EMEE.
Table 4: Estimated marginal excursion effect ($\beta_{0}$) and effect moderation by $\text{Day}_{t}$ (with model $\text{CEE}=\beta_{1}+\beta_{2}\text{Day}_{t}$) of Drink Less push notification on whether the user opens the app in the subsequent 72 hours.
Estimand Method Estimates SE 95% CI $p$-value
pd-EMEE 0.127 0.027 (0.075, 0.180) $<0.001$
$\beta_{0}$ : marginal excursion effect EMEE 0.079 0.056 (-0.031, 0.188) 0.159
pd-EMEE 0.195 0.046 (0.105, 0.286) $<0.001$
$\beta_{1}$: causal excursion effect on the first day EMEE 0.284 0.081 (0.123, 0.445) $<0.001$
pd-EMEE -0.006 0.003 (-0.012, 0.001) 0.092
$\beta_{2}$: change in the causal excursion effect with each additional day EMEE -0.016 0.005 (-0.026, -0.006) 0.002
  • SE, standard error; CI, confidence interval. SE, 95% CI and $p$-value are based on the small-sample correction described in Section 3, Remark 2.
@@ 6 Discussion
We proposed the per-decision estimator of marginal excursion effect (pd-EMEE) for the causal excursion effect of a time-varying treatment on a binary longitudinal outcome for micro-randomized trials. The proposed estimator is applicable as long as the binary outcome encodes the occurrence of an event of interest in a pre-specified time window. When this time window spans multiple decision points, the proposed estimator improves efficiency compared to the currently widely-used EMEE estimator [11], by using a modified form of inverse probability weighting (IPW) called per-decision IPW. This efficiency improvement can lead to more precise estimation of the causal effect or smaller sample size when planning an MRT. The efficiency improvement is greater when there is a higher chance of one or more additional treatments occurring in the proximal outcome window; for example, when the proximal outcome window consists of more decision points or when the randomization probability is larger. When the proximal outcome window doesn’t contain additional decision points (i.e., when the proximal outcome is measured immediately), the proposed method is equivalent to the EMEE in [11].
@@
The method assumes that the binary proximal outcome satisfies a maximum property, i.e., the outcome can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. This holds when the outcome indicates the occurrence of an instantaneous event. For non-instantaneous events, if it is still possible to determine the first decision point, within the proximal outcome window, at which the binary outcome has already occurred, then one can use a similar estimator for efficiency improvement by assuming a generalized maximum property. The generalization is detailed in Supplementary Material A.
@@
The estimator can also be generalized to settings where the number of decision points within a proximal outcome window is not fixed. Such a setting may arise if the decision points are irregularly spaced. In this setting, as long as the occurrence of the event may be determined early on in the proximal outcome window, one can still apply the per-decision IPW idea and thus discard the IPW terms after the event occurs to improve efficiency.
@@
The maximum property may resemble the structure of a discrete-time survival outcome, and may raise the question of whether the proximal outcome should be treated as a survival outcome instead of a binary one. We believe the choice of the model and outcome type should depend on the scientific question. For some interventions, the goal is so that a desirable event, such as self-monitoring in Drink Less and activity bout in HeartStep, occurs within a reasonable time window but not necessarily as soon as possible, and thus the causal effect on the binary outcome contains is relevant for assessing and improving the intervention. For other settings, the goal might be to reduce the risk of a dangerous or harmful behavior as soon as possible, in which case the causal effect on the survival outcome might be more relevant. An example for the latter is problem anger prevention [31], where the intervention is to reduce the risk of aggressive behavior once anger is detected.
@@
We considered the same causal excursion effect as in [11], where the future $\Delta-1$ treatment assignments ($A_{t+1},\ldots,A_{t+\Delta-1}$) are all set to 0. That is, we consider excursions starting from decision point $t$ that receives treatment (or no treatment) at $t$, and then receive no treatment for the next $\Delta-1$ decision points. As pointed out by a reviewer, the choice of this reference treatment regime depends on the scientific question, and in some applications other reference regimes may be of interest. In Supplementary Material H we provide an extension of the proposed method to a class of alternative reference regimes and evaluated the efficiency gain under those settings.
7 Software
R (R Development Core Team [32]) code to reproduce the simulations and the real data analysis can be downloaded at https://github.com/tqian/EMEE_binary_modified_weight. The Drink Less data set is publicly available at https://osf.io/mtcfa.
Acknowledgments
Lauren Bell is supported by a PhD studentship funded by the MRC Network of Hubs for Trials Methodology Research (MR/L004933/2-R18).
@@ Acknowledgments
Appendix A Generalized Version of Maximum Property
@@
$Y_{t,\Delta}=\max(R_{t,t+1},R_{t,t+2},\ldots,R_{t,t+\Delta}).$ (11)
@@
For example, in the HeartSteps II MRT [12], the proximal outcome was whether there is an activity bout in the 30 minutes following a decision point. There was a decision point every 5 minutes, so we denote by $Y_{t,\Delta=6}$ the proximal outcome using our notation. Here $R_{t,t+1}$ is whether an activity bout occurred between $t$ and $t+1$, $R_{t,t+2}$ is whether an activity bout occurred between $t$ and $t+2$, and so on. Here (11) holds with $\Delta=6$.
@@
Appendix B Connection to the Literature
B.1 Connection between Per-Decision IPW (Our Approach) and the Per-Decision Importance Sampling
We describe a related IPW technique in the literature, the per-decision importance sampling (PDIS) in offline reinforcement learning [14], and discuss how our per-decision IPW approach connects to and differs from PDIS. To make connections to our approach, we describe a simple example of PDIS using a setting similar to MRT. See [33, Section 2] for the general form of PDIS. Suppose there are $\Delta$ decision points, and at each decision point a treatment is randomized with randomization probability $p_{t}$ and a short-term outcome $R_{t+1}$ is observed. Let a distal outcome (an outcome of interest) be $Y:=\sum_{t=1}^{\Delta}R_{t+1}$. Suppose the goal is to estimate $E\{Y(a,\bar{0}_{\Delta-1})\}$. Then we have
$E\{Y(a,\bar{0}_{\Delta-1})\}$ $=E\left[\left\{\frac{\mathbbm{1}(A_{1}=a)}{ap_{1}+(1-a)(1-p_{1})}\prod_{s=2}^{\Delta}\frac{\mathbbm{1}(A_{s}=0)}{1-p_{s}}\right\}\sum_{t=1}^{\Delta}R_{t+1}\right]$ (12)
$=E\left[\sum_{t=1}^{\Delta}\left\{\frac{\mathbbm{1}(A_{1}=a)}{ap_{1}+(1-a)(1-p_{1})}\prod_{s=2}^{t}\frac{\mathbbm{1}(A_{s}=0)}{1-p_{s}}\right\}R_{t+1}\right],$ (13)
where (12) is the identification result using the usual IPW and (13) is the identification result using PDIS. The intuition behind (13) is that earlier outcomes cannot depend on later treatments. The same intuition is used in our approach to derive the identification result for proximal outcomes satisfying the maximum property; i.e.,
$=E\left(E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{\mathbbm{1}(A_{j}=0)}{1-p_{j}\left(H_{j}\right)}\right]^{\mathbbm{1}(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\}\mid S_{t},I_{t}=1\right).$ (14)
The difference between our approach and PDIS is that (i) the PDIS literature typically considers the outcome to be a sum of sub-outcomes whereas we considers a binary outcome that is the maximum of sub-outcomes, (ii) the PDIS literature considers a single outcome at the end whereas we considered the setting where the main outcome itself is longitudinal (i.e., we consider $Y_{t,\Delta}$ for $1\leq t\leq T$), and (iii) the PDIS identification result (13) is solely based on IPW, whereas our identification result (14) is a combination of IPW and standardization.
B.2 Connection between Our Estimating Equation and the Literature
Our estimating function is defined as
Besides the per-decision IP weight, $W_{it}$, which was explained in the previous subsection, our estimating function $U_{i}(\alpha,\beta)$ resembles several existing estimating functions in the literature: the multiplicative structural nested mean model (MSNMM) [23], the estimator for causal excursion effect (EMEE) [11], and the weighted and centered least squares (WCLS) estimator [18].
with $\tilde{\epsilon}_{it}(\alpha,\beta)=Y_{it,1}-g(H_{it}^{T}\alpha-(A_{it}-\tilde{p}_{t}(S_{it}))S_{it}^{T}\beta$. Our estimating function (15) adopts the weighting (by $M_{it}$) and the centering of the treatment (by $\tilde{p}_{t}(S_{it})$), which were used to (i) estimate a marginal causal effect instead of a fully conditional causal effect, and (ii) make the estimator robust to misspecified control variables (the $g_{t}(H_{it})$ part).
To extend the WCLS estimating function to binary outcome, the EMEE estimator [11] used a blipping-down technique of MSNMM developed by [23], i.e., pre-multiplying the residual term $\epsilon_{it}(\alpha,\beta)$ by a negative causal effect (to remove the causal effect, i.e., “blip” the outcome down to its counterpart under no treatment). We borrowed this idea from EMEE. This is where the term $e^{-A_{it}S_{it}^{T}\beta}$ comes from. Note that the original blipping-down technique [23] depends on an exponent representing the fully conditional causal effect (such as $e^{-A_{it}f(H_{it})^{T}\theta}$), but we used a blipping-down factor that represents the marginal causal effect because this is the causal effect of interest in our case.
Appendix C Proof of Identifiability Result
@@
Under Assumptions 1-4 in Section 2.4, we have
$E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1\right\}$ (16)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\},$ (17)
which implies the identification result (4) in the main paper. Equation (17) follows immediately from the statements of Lemma C.3 and Lemma C.4 below. We first define a set of new quantities used in the proof, then state and prove Lemmas C.1-C.4.
@@
$Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})\times\mathbbm{1}(C_{tu})=Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1},\bar{a}_{t+u:t+\Delta-1})\times\mathbbm{1}(C_{tu}).$ (18)
Proof 1 (Proof of Lemma C.1)
When $\mathbbm{1}(C_{tu})=0$, (18) holds trivially. When $\mathbbm{1}(C_{tu})=1$, it must be the case that $R_{t+u}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1})=1$, and Assumption 4 implies that $Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1},\bar{a}_{t+u:t+\Delta-1})$ and $Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})$ both equal to 1. Thus (18) holds when the indicator function equals 1. This completes the proof.
Lemma C.2
@@
Proof 2 (Proof of Lemma C.2)
@@
$=$ $E\left\{\prod_{j=t+1}^{t+k_{0}-1}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (19)
$=$ $E\left[E\left\{\prod_{j=t+1}^{t+k_{0}-1}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (20)
@@
×$\left.\times E\left\{\frac{\mathbbm{1}(A_{t+k_{0}}=0)}{P(A_{t+k_{0}}=0\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu})}\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (21)
$=$ $E\left[E\left\{\prod_{j=t+1}^{t+k_{0}}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (22)
$=$ $E\left\{\prod_{j=t+1}^{t+k_{0}}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (23)
Here, (19) follows from the induction assumption, (20) and (23) follow from the law of iterated expectation, (21) follows from the fact that $E\left\{\frac{\mathbbm{1}(A_{t+k_{0}}=0)}{P(A_{t+k_{0}}=0\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu})}\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}=1$ (valid under Assumptions 2 and 3), and (22) follows from Assumption 2 (in particular, conditional on $H_{t+k_{0}}$, $A_{t+k_{0}}$ is independent of all potential outcomes including $Y_{t,\Delta}$ and $C_{tu}$). This completes the proof.
Lemma C.3
@@
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1\right\}$ (24)
@@
Proof 3 (Proof of Lemma C.3)

### v3

@@ 6.1 Drink Less MRT
We applied the proposed pd-EMEE and pd-EMEE2 methods to analyze the Drink Less MRT data. Recall that there are $n=349$ participants in this MRT, each in the study for $T=30$ days. Each participant was randomized once every day at 8 pm with 0.6 probability of receiving a push notification encouraging using the Drink Less app ($A_{t}=1$) and with 0.4 probability of receiving nothing ($A_{t}=0$). Individuals were available to be randomized at all decision points ($I_{t}\equiv 1$). The proximal outcome, $Y_{t,\Delta}$, is whether the participant logged in to the app in the subsequent $\Delta$ days. The binary sub-outcome, $R_{t+1}$, is whether the participant logged in to the app between 8 pm of day $t$ and 7:59 pm of day $t+1$, and $Y_{t,\Delta}=\max(R_{t+1},\ldots,R_{t+\Delta})$. We conducted separate analyses for $\Delta=1,2,\ldots,5.$
[…]
Figure 3: Estimated marginal excursion effect (Moderator: None) and effect moderation by decision point index, past treatment, or past app usage of Drink Less push notification on whether the user opens the app in different subsequent times windows from $\Delta=1$ day to $\Delta=5$ days. For each moderator and estimand, the relative efficiency of pd-EMEE or pd-EMEE2 against EMEE with various $\Delta$ is shown in the corresponding dot figure below.
Figure 3 shows the analysis results, where for each pair of plots the top one shows the point estimate and 95% confidence interval (CI) for the three methods (EMEE, pd-EMEE, pd-EMEE2) and the bottom one shows the estimated relative efficiency ($\widehat{\text{RE}}$) of pd-EMEE or pd-EMEE2 against EMEE. $\widehat{\text{RE}}$ between two estimators is computed as the ratio between the estimated standard error of the estimators. We see that for each estimand and each $\Delta$, EMEE, pd-EMEE, and pd-EMEE2 yield estimates in the same direction with CIs mostly overlapping. Both pd-EMEE and pd-EMEE2 yield substantially narrower CIs than EMEE for $\Delta\geq 2$ and their $\widehat{\text{RE}}$ against EMEE increases with larger $\Delta$ as expected. For example, with $\Delta=3$, the CI for EMEE is almost twice as wide as the CIs for pd-EMEE and pd-EMEE2. pd-EMEE2 is slightly more efficient than pd-EMEE in most cases, especially for the larger $\Delta$ (Δ=4 or $\Delta=4\text{ or }5$).
6.2 HeartSteps I MRT
We applied the proposed pd-EMEE and pd-EMEE2 methods to analyze the HeartSteps I MRT data. Recall that each participant was in the study for 42 days and was randomized five times a day, each time with 0.6 probability of receiving an activity suggestion ($A_{t}=1$) and with 0.4 probability of receiving nothing ($A_{t}=0$). The data set included 37 participants with a total of 7629 decision points, and seven participants had less than 210 decision points of data (with reasons detailed in (Klasnja and others, 2018)). Individuals were available to be randomized ($I_{t}=1$) for about 80% of the decision points, and the reason for unavailability ($I_{t}=0$) includes the person currently driving, having just had an activity bout, or having connection issues. In our analysis, $\Delta=5$ corresponds to a 24-hour proximal outcome window. We considered five different proximal outcomes corresponding to five different step count thresholds. Given a total step count threshold $C$, the proximal outcome, $Y_{t,\Delta}$, is whether the participant takes at least $C$ total steps in the 24-hour window subsequent to decision point $t$ (we omit the notational dependence of $Y_{t,\Delta}$ on $C$). The binary sub-outcome $R_{t,t+s}$ for $1\leq s\leq\Delta$, which is the double-subscript version explicated in the Supplementary Material B, is whether the individual takes at least $C$ total steps between decision points $t$ and $t+s$, and $Y_{t,\Delta}=\max(R_{t,t+1},\ldots,R_{t,t+\Delta})$. We conducted separate analyses for $C=6000,7000,\ldots,10000$.
[…]
Figure 4: Estimated marginal excursion effect (Moderator: None) and effect moderation by decision point index, whether the participant is at home/work, whether the day is weekday of HeartSteps activity suggestion on whether the user takes more than a given threshold of steps in the next 24 hours. For each moderator and estimand, the relative efficiency of pd-EMEE or pd-EMEE2 against EMEE with various step count threshold is shown in the corresponding dot figure below.
Figure 4 shows the analysis results, and the format of the figure is the same as Figure 3. We see that for each estimand and for each threshold $C$, EMEE, pd-EMEE, and pd-EMEE2 yield estimates in the same direction with CIs mostly overlapping. Both pd-EMEE and pd-EMEE2 yield substantially narrower CIs than EMEE for most threshold $C$ values. The efficiency gain of pd-EMEE and pd-EMEE2 against EMEE is not as pronounced as in the Drink Less application but is still substantial, with $\widehat{\text{RE}}$ mostly in the range of 1.2–3.
@@ 7 Discussion
We proposed the per-decision estimator of marginal excursion effect (pd-EMEE) and its projection-based variant (pd-EMEE2) for the causal excursion effect of a time-varying treatment on a binary longitudinal outcome for micro-randomized trials. The proposed estimators are applicable as long as the binary outcome encodes the occurrence of an event of interest in a pre-specified time window. When this time window spans multiple decision points, the proposed estimator improves efficiency compared to the currently widely-used EMEE estimator (Qian and others, 2021), by using a modified form of inverse probability weighting (IPW) called per-decision IPW. The pd-EMEE2, motivated by results from the semiparametric efficiency theory, further improves efficiency. This efficiency improvement can lead to a more precise estimation of the causal effect or a smaller sample size needed when planning an MRT. The efficiency improvement is greater when there is a higher chance of one or more additional treatments occurring in the proximal outcome window; for example, when the proximal outcome window consists of more decision points or when the randomization probability is larger.
@@
The methods require that the binary proximal outcome satisfies a maximum property, i.e., the outcome can be expressed as the maximum of a series of sub-outcomes defined over sub-intervals of time. This holds when the outcome indicates the occurrence of an instantaneous event. For non-instantaneous events, if it is still possible to determine the first decision point, within the proximal outcome window, at which the binary outcome has already occurred, then one can use a similar estimator for efficiency improvement by assuming a generalized maximum property. The generalization is detailed in Supplementary Material B, and estimators based on this generalization are illustrated in Section 6.2.
@@
The estimator can also be generalized to settings where the number of decision points within a proximal outcome window is not fixed. Such a setting may arise if the decision points are irregularly spaced and the proximal outcome window is based on calendar time. In this setting, as long as the occurrence of the event may be determined early on in the proximal outcome window, one can still apply the per-decision IPW idea and thus discard the IPW terms after the event occurs to improve efficiency, with the necessary notational generalization.
@@
The maximum property may resemble the structure of a discrete-time survival outcome, and may raise the question of whether the proximal outcome should be treated as a survival outcome instead of a binary one. We believe the choice of the model and outcome type should depend on the scientific question. For some interventions, the goal is so that a desirable event, such as self-monitoring in Drink Less and activity bout in HeartStep, occurs within a reasonable time window but not necessarily as soon as possible. For such settings, the causal effect on the binary outcome will contain relevant information for assessing and improving the intervention. For other settings, the goal might be to reduce the risk of a dangerous or harmful behavior as soon as possible, in which case the causal effect on the survival outcome might be more relevant. An example of the latter is problem anger prevention (Metcalf and others, 2022), where the intervention is to reduce the risk of aggressive behavior once anger is detected.
@@
We considered the same causal excursion effect as in Qian and others (2021), where the future $\Delta-1$ treatment assignments ($A_{t+1},\ldots,A_{t+\Delta-1}$) are all set to 0. That is, we consider excursions starting from decision point $t$ that receives treatment (or no treatment) at $t$, and then receive no treatment for the next $\Delta-1$ decision points. The choice of this reference treatment regime depends on the scientific question, and in some applications other reference regimes may be of interest. In Supplementary Material J we provide an extension of the proposed method to a class of alternative reference regimes and evaluate the efficiency gain under those settings.
@@
Appendix A Interpretation of the Causal Excursion Effect and Its Use in MRT
The causal excursion effect (2.2) differs from most causal effects under time-varying treatments such as marginal structural models and structural nested mean models (Robins, 1994; Robins and others, 2000). Instead of contrasting fixed treatment trajectories, (2) contrasts between two stochastic treatment trajectories, $(\bar{A}_{t-1},1,\bar{0}_{\Delta-1})$ and $(\bar{A}_{t-1},0,\bar{0}_{\Delta-1})$. That is, (2) is a causal contrast between two excursions from the treatment policy: following the treatment policy up to time $t-1$ and then deviating from the policy to assign $A_{t}=1$ and no treatment for the next $\Delta-1$ decision points, and following the treatment policy up to time $t-1$ and then deviating to assign $A_{t}=0$ and no treatment for the next $\Delta-1$ decision points. (2) is conditional on $I_{t}(\bar{A}_{t-1})=1$ and $S_{t}(\bar{A}_{t-1})$, meaning that it concerns the subpopulation who, after following the treatment policy up to time $t-1$, are available to be randomized at $t$ and within a stratum defined by $S_{t}(\bar{A}_{t-1})$. In practice we often impose a model on how $\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ depends on $S_{t}(\bar{A}_{t-1})$ to pool across $S_{t}$-strata.
$\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ depends on the treatment policy in the MRT, because the variables in $H_{t}(\bar{A}_{t-1})\setminus S_{t}(\bar{A}_{t-1})$ are marginalized over in the conditional expectations in (2). Such dependence is scientifically desirable for two reasons. First, the causal excursion effect approximates the treatment effect in a real-world implementation. A good MRT policy would already have incorporated implementation considerations such as burden and feasibility through the choice of randomization probability and availability criteria, so a feasible policy will not deviate too far from the MRT policy. Second, the causal excursion effect indicates the effective deviations from the MRT policy and how the policy might be improved. A fully marginal effect (by setting $S_{t}(\bar{A}_{t-1})=1$ in (2)) indicates whether a treatment is worth further investigation, and an effect modification analysis (by setting $S_{t}(\bar{A}_{t-1})$ to be certain time-varying covariates in (2)) indicates whether the MRT policy should be modified to depend on time-varying covariates. In addition, the dependence of $\text{CEE}\left\{S_{t}(\bar{A}_{t-1})\right\}$ on the MRT policy resembles the primary analysis in factorial designs and allows one to design trials with a higher power to detect meaningful effects. Related marginalization ideas were considered by Robins (2004); Neugebauer and others (2007). See Boruvka and others (2018), Qian and others (2021), and Qian and others (2022) for more discussion on causal excursion effect and its use in MRT. See Guo and others (2021) for a comprehensive comparison of various causal estimands under time-varying treatments.
Appendix B Generalized Version of Maximum Property
@@
$Y_{t,\Delta}=\max(R_{t,t+1},R_{t,t+2},\ldots,R_{t,t+\Delta}).$ (B.1)
@@
For example, in the HeartSteps II MRT (Liao and others, 2020), the proximal outcome was whether there is an activity bout in the 30 minutes following a decision point. There was a decision point every 5 minutes, so we denote by $Y_{t,\Delta=6}$ the proximal outcome using our notation. Here $R_{t,t+1}$ is whether an activity bout occurred between $t$ and $t+1$, $R_{t,t+2}$ is whether an activity bout occurred between $t$ and $t+2$, and so on. Here (B.1) holds with $\Delta=6$.
@@
Appendix C Causal Assumptions and Parameter Identification
@@
Under Assumptions 1-4, we have
$E\left\{Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1\right\}$ (C.1)
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}=0)}Y_{t,\Delta}\mid A_{t}=a,H_{t},I_{t}=1\right\},$ (C.2)
which implies the identification result (4) in the main paper. Equation (C.2) follows immediately from the statements of Lemma 5 and Lemma 6 below. We first define a set of new quantities used in the proof, then state and prove Lemmas 3-6.
@@
$Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})\times\mathbbm{1}(C_{tu})=Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1},\bar{a}_{t+u:t+\Delta-1})\times\mathbbm{1}(C_{tu}).$ (C.3)
Proof 1 (Proof of Lemma 3)
When $\mathbbm{1}(C_{tu})=0$, (C.3) holds trivially. When $\mathbbm{1}(C_{tu})=1$, it must be the case that $R_{t+u}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1})=1$, and Assumption 4 implies that $Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{u-1},\bar{a}_{t+u:t+\Delta-1})$ and $Y_{t,\Delta}(\bar{A}_{t-1},a_{t},\bar{0}_{\Delta-1})$ both equal to 1. Thus (C.3) holds when the indicator function equals 1. This completes the proof.
Lemma 4
@@
Proof 2 (Proof of Lemma 4)
@@
$=$ $E\left\{\prod_{j=t+1}^{t+k_{0}-1}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}$ (C.4)
$=$ $E\left[E\left\{\prod_{j=t+1}^{t+k_{0}-1}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (C.5)
@@
×$\left.\times E\left\{\frac{\mathbbm{1}(A_{t+k_{0}}=0)}{P(A_{t+k_{0}}=0\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu})}\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (C.6)
$=$ $E\left[E\left\{\prod_{j=t+1}^{t+k_{0}}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right]$ (C.7)
$=$ $E\left\{\prod_{j=t+1}^{t+k_{0}}\frac{\mathbbm{1}(A_{j}=0)}{p(A_{j}=0\mid H_{j})}\times Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1,C_{tu}\right\}.$ (C.8)
Here, (C.4) follows from the induction assumption, (C.5) and (C.8) follow from the law of iterated expectation, (C.6) follows from the fact that $E\left\{\frac{\mathbbm{1}(A_{t+k_{0}}=0)}{P(A_{t+k_{0}}=0\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu})}\mid A_{t}=a,H_{t+k_{0}},I_{t}=1,C_{tu}\right\}=1$ (valid under Assumptions 2 and 3), and (C.7) follows from Assumption 2 (in particular, conditional on $H_{t+k_{0}}$, $A_{t+k_{0}}$ is independent of all potential outcomes including $Y_{t,\Delta}$ and $C_{tu}$). This completes the proof.
Lemma 5
@@
$=$ $E\left\{\prod_{j=t+1}^{t+\Delta-1}\left[\frac{1(A_{j}=0)}{P(A_{j}=0\mid H_{j})}\right]^{1(\max_{1\leq s\leq j-t}R_{t+s}(\bar{A}_{t-1},a,\bar{0})=0)}Y_{t,\Delta}(\bar{A}_{t-1},a,\bar{0})\mid A_{t}=a,H_{t},I_{t}=1\right\}$ (C.9)
@@
Proof 3 (Proof of Lemma 5)